Circular No. 6876

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SAX J1808.4–3658 = XTE J1808–369 F. E. Marshall, Goddard Space Flight Center, on behalf of the RXTE

Science Operations Center, reports the detection of a transient x-ray source with the Proportional Counter Array on the Rosssi X-ray Timing Explorer at α = 18^{h}08^{m}34^{s} ± 4^{s} (90-percent confidence interval), δ = −36^{o}59′.4 ± 3′ (equinox 2000.0): “The 2–10-keV flux was 8 × 10^{−10} erg cm^{−2} on Apr. 9.6 UT and increased to 1.5 × 10^{−9} erg cm^{−2} on Apr. 11.8. The position is consistent with the x-ray transient SAX J1808.4–3658 (in ’t Zand et al. 1998, A.Ap. 331, L25), for which type-I x-ray bursts were seen. A continuing x-ray lightcurve from the RXTE All Sky Monitor is available

at http://heasarc.gsfc.nasa.gov/xte weather. Observations at other

wavelengths are encouraged.”

R. Wijnands and M. van der Klis, Astronomical Institute ‘Anton Pannekoek’, University of Amsterdam, report the discovery of near-coherent pulsations between 400.98 and 401.02 Hz (2.49 ms) in the persistent emission of the x-ray transient XTE J1808–369, in RXTE PCA observations taken between Apr. 11.32 and 11.37 UT: “An approximately sinusoidal variation is observed in the frequency of these oscillations, part of which is due to the Doppler shifts caused by the satellite motion. The strength of the oscillations was 4.2 percent rms in the photon energy range 2–60 keV, with little dependence on photon energy. The 2–60-keV flux was ∼ 60 mCrab. The presence of the pulsations was confirmed in data taken on Apr. 13 using a different RXTE data mode. This is the first x-ray source showing nearly coherent millisecond oscillations in its persistent emission. The most likely interpretation is the presence of an accretion-powered millisecond x-ray pulsar in the system. XTE J1808–369 has a position that is consistent with that of the known x-ray transient and x-ray burster SAX J1808.4–3658. The best BeppoSAX coordinates are α = 18^{h}08^{m}29^{s}, δ = −36^{o}58′.6 (equinox 2000.0; error radius 2′; in ’t Zand et al., op.cit.). More x-ray observations are needed in order to study this unique system in more detail. Observations at other wavelengths are encouraged in order to determine the exact nature of this system.”

SUPERNOVA 1998aq IN NGC 3982 Magnitude obtained by D. Hanzl, Brno, Czech Republic (0.4-m reflec-

tor + CCD + V filter; GSC comparison stars): Apr. 14.834 UT, 14.3.

1998 April 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green